Ana is a 15-year-old high school student who is active in school activities, loves to run and bike, and has a 3.95 grade poinCalculate BMI and percentile from the sports physical visit.

a. 17.0, 9th percentile

b. 19.4, 20th percentile

c. 16.1, 6th percentile

d. 17.0, 15th percentile

How would you classify her weight status at the sports physical visit?

a. Ana's weight at the sports physical visit is underweight according to the CDC

b. Ana's weight at the sports physical is still healthy according to the CDC

c. Ana's weight at the sports physical is severely underweight according to the CDC

d. Ana's weight at the sports physical is at the high end of normal according to the CDC
Answer #1

BMI=Weight  in kg divided by height in meter square

BMI=kg÷m2

Here Ana's weight=96pounds (43.54kg)

Height=5'3"(1.6m)

BMI =43.54÷(1.6)2

=43.54÷2.56

=17.0

In accordance to her BMI she falls in 15th percentile which is a healthy weight(normal range 16.8 to 23.2 in the chart )

The percentile range are

  • It is considered underweight, if it lies between 0 to 4
  • It is healthy weight, if it falls between 5 to 85
  • It is over weight, if it falls between 86 to 95
  • It is obesity ,if it falls above 96

Ans:d. 17.0,15th percentile

The BMI of 17 year old girl at 15 years is normal and it indicates a healthy weight

Ans:a.Ana's weight at the sports physical is still healthy according to the CDC